# Leveraging Network Methods for Hub-like Microservice Detection

### v1

@@ I Introduction
Microservice Architecture (MSA) is a popular software development paradigm for scalable and manageable systems [1], which facilitates flexibility by decomposing applications into small, independently deployable services. However, with this architecture comes its challenges that can result in design flaws that negatively impact system quality. Catalogs of architectural anti-patterns have been proposed [2] to highlight the negative aspects of flawed microservice design. Detection of these anti-patterns is crucial in ensuring the quality of microservice-based systems [3].
@@ III Related Work
Other aspects of the network are not considered. Different aspects of the MSA network can be considered to determine whether the service is Hub-like or not. Section II presented several definitions of centrality metrics that take different topological properties into account, like consolidating paths or pointing to influential nodes. Works that do discuss and compute other centralities ([30, 31, 32]) do not apply them to detect a particular pattern but rather just consider the service ranking w.r.t the computed centralities, labeling the most central services as potentially problematic (Table I). Bakhtin et al. [5] recently presented network centrality as a new perspective on microservice architecture and proposed it to be used for anti-pattern detection. Kirkley [6] on the other hand, argues against using degree, degree centrality or any other centrality-based ranking, since it requires either setting a threshold or defining the notion of too high/too many, usually through an outlier detection method, which comes with its own set of parameters. Instead, he argues for developing completely non-parametric methods for hub detection.
@@ IV-B Study Context
We used the dataset of microservice architectural networks provided by Bakhtin et al. [5]. The authors compiled a dataset of OSS MSS projects suitable for network analysis by reconstructing their SDGs. They leveraged Code2DFD tool [39] to reconstruct the architecture statically, which only applies to the Java Spring framework. They thus identified 125 Java Spring systems from four datasets of OSS MSS [40, 41, 42, 43]. These 125 Java Spring projects provide complete architectures of 24 OSS MSS projects reconstructed successfully with the Code2DFD tool. We took the 24 SDGs directly from [5]. Furthermore, we used two versions of the train-ticket benchmark project to enrich our data since the architecture varies noticeably between releases v0.0.1 and v1.0.0 of the project, and thus can provide additional data points for analysis. We followed the same procedure as Bakhtin et al. [5] to obtain the network of train-ticket v0.0.1. The list of the analyzed projects in available in the Replication Package [44].
@@ VI Discussion
The implication of our results for practitioners is thus that there is reasonable evidence to consider utilizing either the normalized amount of connections of the microservice, i.e., the degree centrality, as the guiding metric for reasoning about the microservices’ connections and detecting the Hub-like anti-pattern, or leveraging the ER hub detection method by Kirkley [6] in the Arcan or similar tool, since these approaches enable more robust and generalizable analysis.

### v2

@@ I Introduction
Microservice Architecture (MSA) is a popular software development paradigm for scalable and manageable systems [1], which facilitates flexibility by decomposing applications into small, independently deployable services. However, with this architecture comes its challenges that can result in design flaws that negatively impact system quality. Catalogs of architectural anti-patterns have been proposed [2, 3] to highlight the negative aspects of flawed microservice design [4]. Detection of these anti-patterns is crucial in ensuring the quality of microservice-based systems [5].
@@ III Related Work
Other aspects of the network are not considered. Different aspects of the MSA network can be considered to determine whether the service is Hub-like or not. Section II presented several definitions of centrality metrics that take different topological properties into account, like consolidating paths or pointing to influential nodes. Works that do discuss and compute other centralities ([32, 33, 34]) do not apply them to detect a particular pattern but rather just consider the service ranking w.r.t the computed centralities, labeling the most central services as potentially problematic (Table I). Bakhtin et al. [7, 39] recently presented network centrality as a new perspective on microservice architecture and proposed it to be used for anti-pattern detection. Kirkley [8] on the other hand, argues against using degree, degree centrality or any other centrality-based ranking, since it requires either setting a threshold or defining the notion of too high/too many [40], usually through an outlier detection method [41], which comes with its own set of parameters. Instead, he argues for developing completely non-parametric methods for hub detection.
@@ IV-B Study Context
We used the dataset of microservice architectural networks provided by Bakhtin et al. [7]. The authors compiled a dataset of OSS MSS projects suitable for network analysis by reconstructing their SDGs. They leveraged Code2DFD tool [44, 45] to reconstruct the architecture statically, which only applies to the Java Spring framework. They thus identified 125 Java Spring systems from four datasets of OSS MSS [46, 47, 48, 49]. These 125 Java Spring projects provide complete architectures of 24 OSS MSS projects reconstructed successfully with the Code2DFD tool. We took the 24 SDGs directly from [7]. Moreover, we took an additional version of the train-ticket benchmark project to enrich our data from [39] since the architecture varies noticeably between releases v0.0.1 and v1.0.0 of the project, and thus can provide additional data points for analysis. The list of the analyzed projects in available in the Replication Package [50].
@@ VI Discussion
The implication of our results for practitioners is thus that there is reasonable evidence to consider utilizing either the normalized amount of connections of the microservice, i.e., the degree centrality, as the guiding metric for detecting the Hub-like anti-pattern, or leveraging the ER hub detection method by Kirkley [8] in the Arcan or similar tool, since these approaches enable more robust and generalizable analysis.